OpenStudy (anonymous):

A quilter made 256 small squares for a large quilt. If the quilt is shaded like a square, how many small squares will she use on each side

OpenStudy (dumbcow):

256 = 16^2

OpenStudy (anonymous):

Shouldn't there be two sides? in that case 256 = 2 *128 11^2 = 121, so on each side 121 2*121 = 242 and 14 left over Unless I misinterprete the question